OpenStudy (anonymous):

Suppose w=a+bi is one of the 31st roots of 1. a. What is the maximum value of a? b. What is the minimum value of b? Please help! What does this even mean?!? >.<

OpenStudy (anonymous):

@ganeshie8 @mathslover anyone?

OpenStudy (anonymous):

Do you know how to find the complex roots of a number, to begin with?

OpenStudy (anonymous):

Like using De Moivre's Theorem?

ganeshie8 (ganeshie8):

finding max value of a+bi is easy : since "1" itself is one of the 31st roots of 1, max value for a = 1

ganeshie8 (ganeshie8):

you can find the min value for "b" using De Moivre's thm : 31st roots = \(\large e^{i \frac{2\pi k}{31}}\)

ganeshie8 (ganeshie8):

\(k = 0, 1, ... 30\)

ganeshie8 (ganeshie8):



ganeshie8 (ganeshie8):

test those two points ^

ganeshie8 (ganeshie8):

31 * 3/4 = 23.25 so the min value occurs at k = 23 or 24

OpenStudy (anonymous):

Hmmm I understand the max value part, but my answer key says the min value should be at \[\sin(16\pi/31)\] or about 0.9987

ganeshie8 (ganeshie8):

\(\large e^{i \frac{2\pi *23}{31}}\) is clearly one of the 31st roots of 1, \(\large e^{i \frac{2\pi *23}{31}} = \cos (\frac{46 \pi}{31}) + i \sin (\frac{46 \pi}{31})\) \(\implies b = \sin (\frac{46 \pi}{31})\)

ganeshie8 (ganeshie8):

\(\sin (\frac{46 \pi}{31}) < \sin (\frac{16 \pi}{31})\) so the min value is \(\sin (\frac{46 \pi}{31})\)

ganeshie8 (ganeshie8):

^^thats for minimum value of \(b\) , the imaginary part

ganeshie8 (ganeshie8):

i think your text book is talking about min value of \(a\) ?

OpenStudy (anonymous):

Oh my fault, it was asking for the MAXIMUM of b, not the minimum. Sorry :S

ganeshie8 (ganeshie8):

oh ok then it makes sense :)

ganeshie8 (ganeshie8):

maximum for imaginary part occurs at TOP, 1/4th the way to full revolution : k = 31*1/4 = 7.75

ganeshie8 (ganeshie8):

so, test k = 7 and k = 8 points

ganeshie8 (ganeshie8):